# Learning Exactly Linearizable Deep Dynamics Models

### v1

@@ 3 Control Design
In this section, we describe a control method that satisfies constraints (1) and (2) and tracks the output $y$ to the reference value $y_{\mathrm{d}}$. This constrained reference-tracking control can be designed as a convex MPC using the S–HW model; however, the introduction of the output feedback in (10) does not ensure the convexity of the input admissible set in the finite-horizon optimal control problem. In this paper, we propose an alternative design method for an EL model that avoids this problem and attributes the control law to convex optimization instead.
3.1 Tracking Control
From the bijectivity of $\Phi$ and $\Psi$ in Definition 1, the tracking control problem for $y$ (using input $v$) can be reduced to a linear control problem for $x$ (using input $u$) for which the optimal design methods are well developed. For the simplicity of discussion, disturbance $d$ is assumed to be constant in the transformed linear dynamics in (10)–(12), such that
$\dot{x}$ $=A(\bar{d})x+B(\bar{d})u+c(\bar{d}),$
$x$ $=\Phi(y,\bar{d}),\ u=\Psi^{-1}(v,y,\bar{d}).$
Suppose that $y_{\mathrm{d}}$ is realizable as a steady state with $d=\bar{d}$; then, there exists $u_{\mathrm{d}}$ such that
$A(\bar{d})x_{\mathrm{d}}+B(\bar{d})u_{\mathrm{d}}+c(\bar{d})=0,\ x_{\mathrm{d}}:=\Phi(y_{\mathrm{d}},\bar{d}).$ (17)
We consider the control input $u(t)$ for the linear part that minimizes the cost function
$J:=\int_{0}^{\infty}\tilde{x}^{\mathrm{T}}Q\tilde{x}+\tilde{u}^{\mathrm{T}}R\tilde{u}\ \mathrm{d}t$ (18)
where $\tilde{x}:=x-x_{\mathrm{d}},\ \tilde{u}:=u-u_{\mathrm{d}}$. The optimal $u$ is given by a linear quadratic regulator
$u=u_{\mathrm{d}}-K\tilde{x},\ K\coloneqq R^{-1}B^{\mathrm{T}}P$ (19)
@@ 3.1 Tracking Control
The control input $v(t)$ is obtained by performing an inverse transformation of (10) for the control input $u(t)$ in (19). The overall flow of the calculation of the control input $v(t)$ is shown in Fig. 3. The controller calculates the control input $v(t)$ from the output signal $y(t)$ obtained from the system to be controlled and the desired target value $y_{\mathrm{d}}$. In the controller, the output signal $y(t)$ is converted into the state variable $x(t)$ through the bijective map $\Phi$, and the internal input signal $u(t)$, which is linearly related to $x(t)$, is obtained using the linear control method (19). In addition, the control input $v(t)$ is obtained by mapping the internal input signal $u(t)$ using the bijective map $\Psi$.
Figure 2: Controller implementation
@@
Fig. 3 shows the resulting closed-loop system comprising Figs. 1 and 3, which perform feedback control on the linear dynamics $\dot{x}=Ax+Bu+c$. The structure is only a linear system in which the bijective maps $\Psi^{-1}$ and $\Phi^{-1}$ on both sides disappear. Figs. 3 and 3 omit the disturbance $\bar{d}$. The nonlinear optimal control theory can be employed for a more advanced control law than the linear feedback described here; see Appendix 2.
@@ 3.2 Constraint-Aware Control
Thus far, constraints (1) and (2) have not been considered. We introduce the control barrier function as one of the methods to satisfy these constraints. Let us consider an input-affine nonlinear system
$\dot{x}$ $=f(x)+g(x)u,$ (21)
$z^{\prime}$ $=h(x)$ (22)
with $x\in X(\subset\mathbb{R}^{n}),u\in U(\subset\mathbb{R}^{m}),z^{\prime}\in\mathbb{R},h:X\rightarrow\mathbb{R}$. Note that (21) corresponds to (9) when applied to our EL models, and $U$ corresponds to the input constraint (1). A continuous and differentiable function $h:X\rightarrow\mathbb{R}$ is called a control barrier function (CBF) if a class $\mathcal{K}_{\infty}$ function is 111 Function $\alpha$ is said to be of class $\mathcal{K}_{\infty}$ if $\alpha$ is a strictly monotonically increasing function, and $\alpha(0)=0,\alpha(r)\rightarrow\infty(r\rightarrow\infty)$. $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ such that $\sup_{u\in U}\left\{\mathcal{L}_{f}h(x)+\mathcal{L}_{g}h(x)u\right\}\leq\alpha(-h(x))$ for any $x\in X$, where $\mathcal{L}_{f}h(x):=\frac{\partial h}{\partial x}f(x)$. We can verify that if $x(0)\in\{x\in X:h(x)\leq 0\}=:\mathcal{C}$ and $u(t)$ is chosen such as $\mathcal{L}_{f}h(x)+\mathcal{L}_{g}h(x)u\leq\alpha(-h(x))$ for any $t$, then $x(t)$ does not exit $\mathcal{C}$. To employ this property for the required constraint satisfaction, we can use the solution $u^{*}$ for the quadratic optimization
$\begin{cases}u^{\ast}=&\mathop{\mathrm{arg~{}min}}\limits_{u\in U}\|u-k(x)\|^{2}\\ \mathrm{s.t.}&\mathcal{L}_{f}h(x)+\mathcal{L}_{g}h(x)u\leq\alpha(-h(x))\end{cases}$ (23)
as a control input, where $k(x)$ is a desirable control law, such as LQR (19). Thus, the output constraint $z^{\prime}=h(x)\leq 0$ is satisfied. In the above description, the constrained output $z^{\prime}$ is assumed to be scalar for simplicity; however, it can be easily extended to a vector by replicating the function $h$. We consider setting $h(x,u)=\Xi(x,u)-\bar{z}=:h_{z}(x,u)$ to handle output constraint (2). However, the aforementioned framework is not directly applicable to this function because of the input dependency of $\Xi$, and the above discussion is only valid if $h$ depends only on $x$. To circumvent this issue, we restrict input signals to continuously differentiable ones, such as
$\dot{u}=\lambda,$ (24)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}x\\ u\\ \end{pmatrix}$ $=\begin{pmatrix}f(x)+g(x)u\\ \lambda\\ \end{pmatrix},$ (29)
$z^{\prime}$ $=h(x,u).$ (30)
Considering an augmented vector $[x^{\mathrm{T}},u^{\mathrm{T}}]^{\mathrm{T}}$ as a state and $\lambda$ as a new input, this system falls into the standard CBF framework because the input dependence of $h$ can be eliminated. The CBF obtained by this augmentation is called the integral CBF (I-CBF) (Ames et al., 2021). From a different perspective, the adoption of I-CBF allows the model to have an input dependence on the function $\Xi$. This motivated us to solve this problem
$\begin{cases}\lambda^{\ast}=&\mathop{\mathrm{arg~{}min}}\limits_{\lambda}\frac{\mathrm{d}}{\mathrm{d}t}\|u-k(x)\|^{2}+a\|\lambda\|^{2}\\ \mathrm{s.t.}&\frac{\partial h}{\partial x}(f(x)+g(x)u)+\frac{\partial h}{\partial u}\lambda\leq\alpha(-h(x,u))\end{cases}$ (31)
with a positive constant $a$. The first term aims at the quick convergence of $u$ to $k(x)$, and the second term is added to regularize the derivative of the input $u$. Although the input constraint $u\in U$ originally considered in (23) disappears in the above problem, it can be treated using this I-CBF approach. We can combine the vector function
$h_{v}(x,u):=\begin{bmatrix}u-\Psi^{-1}(\bar{v},\Phi^{-1}(x))\\ \Psi^{-1}(\underline{v},\Phi^{-1}(x))-u\end{bmatrix}$ (34)
to $h(x,u)$ for handling the input constraint (1). In summary, this approach can handle the constraints (1) and (2) by setting $h(x,u)=[h_{z}(x,u)^{\mathrm{T}},h_{v}(x,u)^{\mathrm{T}}]^{\mathrm{T}}$. The function to be minimized in (31) is quadratic with respect to $\lambda$, and it is given by
$F(\lambda)$ $:=\frac{\mathrm{d}}{\mathrm{d}t}\|u-k(x)\|^{2}+a\|\lambda\|^{2}$ (35)
$=a\|\lambda\|^{2}+2(u-k(x))^{\mathrm{T}}\lambda-2(u-k(x))^{\mathrm{T}}\frac{\partial k(x)}{\partial x}(f(x)+g(x)u).$
In addition, any constraint expressed in the form $h(x,u)\leq 0$ is reduced to the affine constraint for $\lambda$ in the I-CBF approach. Thus, the above optimization problem is QP and can be solved quickly using many ready-made solvers. Further, the EL model allows functions $h(x,u)$ corresponding to (1) and (2) to be convex to $u$ for any fixed $x$. This guarantees that $x,u$ at the equilibrium point of this system is the solution to the problem $\mathop{\mathrm{minimize}}\limits_{u}\|u-k(x)\|^{2}\ \mathrm{s.t.}\ h(x,u)\leq 0$ (See Appendix 3).
@@ 4 Experimental Evaluation
  • Reference $y_{\mathrm{d}}$ for the output $y$ (black solid),
  • Bounds for the input $v$ and output $z$ (black dashed),
@@
Therefore, QP realized the tracking control of output $y$ to the reference $y_{\mathrm{d}}$ while satisfying constraints (1) and (2). For input $v$, the control law based on (19) (blue line) fails to satisfy the constraint (1) (black dashed line), whereas the control law based on QP (green line) is within the constraints. For output $y$, the results of the QP deviate slightly from the reference $y_{\mathrm{d}}$ because there is no input $v$ such that $u=\Psi^{-1}(v,y_{\mathrm{d}},\bar{d})$ satisfies constraints (1) and (2). In other words, from (17), it is possible to obtain $x_{\mathrm{d}}$ and $u_{\mathrm{d}}$ corresponding to $y_{\mathrm{d}}$; however, the $v,z$ obtained by $\Psi$ in (10) and $\Xi$ in (13) do not satisfy the constraints (1), (2). Fig. 7 suggests that our proposed method automatically find a suitable steady input $v$ that achieves good tracking performance for the target $y_{\mathrm{d}}$ while satisfying constraints (1) and (2), which is the purpose of this study.
Figure 5: Results of model prediction
Table 1: Coefficient of determination $R^{2}$
Figure 6: Control result of input $v$
output model S–HW EL(A) EL(B)
$y_{1}$ 0.968 0.974 0.976
$y_{2}$ 0.841 0.905 0.948
$y_{3}$ 0.802 0.960 0.956
$z_{1}$ 0.883 0.890 0.925
$z_{2}$ 0.947 0.966 0.970
Average 0.888 0.939 0.955
Figure 5: Results of model prediction
Table 1: Coefficient of determination $R^{2}$
Figure 6: Control result of input $v$
Figure 7: Control result of output $y,z$

### v2

@@ 3 Control Design
In this section, we describe a control method that satisfies constraints (1) and (2) and regulates the output $x$ to the reference value $x_{\mathrm{d}}$. This constrained regulation control can be designed as a convex MPC using the S–HW model as shown by Moriyasu et al. (2022); however, the introduction of the output feedback in (10) does not ensure the convexity of the input admissible set in the finite-horizon optimal control problem. In this paper, we propose an alternative design method for an EL model that avoids this problem and attributes the control law to convex optimization instead.
3.1 Regulation Control
From the bijectivity of $\Phi$ and $\Psi$ in Definition 1, the regulation control problem for $x$ (using input $u$) can be reduced to a linear control problem for $\xi$ (using input $v$) for which the optimal design methods are well developed. For the simplicity of discussion, exogenous input $d$ is assumed to be constant in the transformed linear dynamics in (10)–(12), such that
$\dot{\xi}$ $=A(\bar{d})\xi+B(\bar{d})v+c(\bar{d}),$
$\xi$ $=\Phi(x,\bar{d}),\ v=\Psi(u,x,\bar{d}).$
Suppose that $x_{\mathrm{d}}$ is realizable as a steady state with $d=\bar{d}$ and there exists $v_{\mathrm{d}}$ such that
$A(\bar{d})\xi_{\mathrm{d}}+B(\bar{d})v_{\mathrm{d}}+c(\bar{d})=0,\ \xi_{\mathrm{d}}:=\Phi(x_{\mathrm{d}},\bar{d}).$ (24)
We consider the control input $v(t)$ for the linear part that minimizes the cost function
$J:=\int_{0}^{\infty}\tilde{\xi}^{\mathrm{T}}Q\tilde{\xi}+\tilde{v}^{\mathrm{T}}R\tilde{v}\ \mathrm{d}t$ (25)
where $\tilde{\xi}:=\xi-\xi_{\mathrm{d}},\ \tilde{v}:=v-v_{\mathrm{d}}$. The optimal $v$ is given by a linear quadratic regulator
$v=v_{\mathrm{d}}+K\tilde{\xi},\ K\coloneqq-R^{-1}B^{\mathrm{T}}P$ (26)
@@ 3.1 Regulation Control
The control input $u(t)$ is obtained by performing an inverse transformation of (10) for the control input $v(t)$ in (26).
Since $\xi$ and $v$ have no physical meaning, determining $Q$ and $R$ to achieve desirable performance is generally difficult. Therefore, it is desirable to be able to consider another objective function $J_{\rm r}=\int_{0}^{\infty}\tilde{x}^{\mathrm{T}}Q_{\rm r}\tilde{x}+\tilde{u}^{\mathrm{T}}R_{\rm r}\tilde{u}\ {\rm d}t$, where $\tilde{x}:=x-x_{\rm d},\tilde{u}:=u-u_{\rm d},u_{\rm d}:=\Psi^{-1}(v_{\rm d},x_{\rm d},\bar{d})$. This function is difficult to minimize directly, but for example, linearization $\tilde{x}\approx(\partial\Phi/\partial x)|_{(x_{\rm d},d)}^{-1}\tilde{\xi},\tilde{u}\approx(\partial\Psi/\partial u)|_{(u_{\rm d},x_{\rm d},d)}^{-1}\tilde{v}$ yields $J_{\rm r}\approx J$ around the target point $x_{\rm d},u_{\rm d}$ with $Q=(\partial\Phi/\partial x)^{\mathrm{T}}Q_{\rm r}(\partial\Phi/\partial x),R=(\partial\Psi/\partial u)^{\mathrm{T}}R_{\rm r}(\partial\Psi/\partial u)$. This can assist designing parameters $Q,R$.
The overall flow of the calculation of the control input $u(t)$ is shown in Fig. 4. The controller calculates the control input $u(t)$ from the output signal $x(t)$ obtained from the system to be controlled and the desired target value $x_{\mathrm{d}},u_{\mathrm{d}}$ satisfying (24) and $u_{\rm d}=\Psi^{-1}(v_{\rm d},x_{\rm d},\bar{d})$. In the controller, the output signal $x(t)$ is converted into the state variable $\xi(t)$ through the bijective map $\Phi$, and the internal input signal $v(t)$, which is linearly related to $\xi(t)$, is obtained using the linear control method (26). In addition, the control input $u(t)$ is obtained by mapping the internal input signal $v(t)$ using the bijective map $\Psi^{-1}$. Fig. 4 shows the resulting closed-loop system comprising Figs. 1 and 4, which perform feedback control on the linear dynamics $\dot{\xi}=A\xi+Bv+c$. The structure is only a linear system in which the bijective maps $\Psi^{-1}$ and $\Phi$ on both sides disappear.
The above design is based on unconstrained linear-quadratic regulator (LQR) and is one of the simplest methods that takes advantage of the internal linearity of the EL model. For a more advanced control, nonlinear optimal control theory can also be employed (see Appendix 2).
@@ 3.2 Constraint-Aware Control
Thus far, constraints (1) and (2) have not been considered. We employ the control barrier function (CBF) as one of the methods to satisfy these constraints.
First, we preliminarily introduce the standard CBF approach considering an input-affine single-output nonlinear system:
$\dot{\xi}$ $=f(\xi)+g(\xi)v,$ (28)
$y^{\prime}$ $=h(\xi)$ (29)
with $\xi\in X(\subset\mathbb{R}^{n}),v\in V(\subset\mathbb{R}^{m}),y^{\prime}\in\mathbb{R},h:X\rightarrow\mathbb{R}$. Note that (28) corresponds to (9) when applied to our EL models, and $V$ corresponds to the input constraint (1). A continuous and differentiable function $h:X\rightarrow\mathbb{R}$ is called a CBF if a class $\mathcal{K}_{\infty}$ function [^1] Function $\alpha$ is said to be of class $\mathcal{K}_{\infty}$ if $\alpha$ is a strictly monotonically increasing function, and $\alpha(0)=0,\alpha(r)\rightarrow\infty(r\rightarrow\infty)$. $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ exists such that $\inf_{v\in V}\left\{\mathcal{L}_{f}h(\xi)+\mathcal{L}_{g}h(\xi)v\right\}\leq\alpha(-h(\xi))$ holds for any $\xi\in X$, where $\mathcal{L}_{f}h(\xi):=\frac{\partial h}{\partial\xi}f(\xi)$. We can verify that if $\xi(0)\in\{\xi\in X:h(\xi)\leq 0\}=:\mathcal{C}$ and $v(t)$ is chosen such that $\mathcal{L}_{f}h(\xi)+\mathcal{L}_{g}h(\xi)v\leq\alpha(-h(\xi))$ holds for any $t$, then $\xi(t)$ does not exit $\mathcal{C}$. To employ this property for the required constraint satisfaction, we can use the solution $v^{*}$ for the quadratic optimization
$\begin{cases}v^{\ast}=&\mathop{\mathrm{arg~{}min}}\limits_{v\in V}\|v-k(\xi)\|^{2}\\ \mathrm{s.t.}&\mathcal{L}_{f}h(\xi)+\mathcal{L}_{g}h(\xi)v\leq\alpha(-h(\xi))\end{cases}$ (30)
as a control input, where $k(\xi)$ is a desirable control law, such as LQR (26). Thus, the output constraint $y^{\prime}=h(\xi)\leq 0$ is satisfied. In the above description, the constrained output $y^{\prime}$ is assumed to be scalar for simplicity; however, it can be easily extended to a vector by replicating the function $h$.
Next, we consider setting $h(\xi,v)=\Xi(\xi,v)-\bar{y}=:h_{y}(\xi,v)$ to handle output constraint (2). However, the aforementioned framework is not directly applicable to this function because of the input dependency of $\Xi$, and the above discussion is only valid if $h$ depends only on $\xi$. To circumvent this issue, we restrict input signals to continuously differentiable ones, such as
$\dot{v}=\lambda,$ (31)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}\xi\\ v\\ \end{pmatrix}$ $=\begin{pmatrix}f(\xi)+g(\xi)v\\ \lambda\\ \end{pmatrix},$ (32)
$y^{\prime}$ $=h(\xi,v).$ (33)
Considering an augmented vector $[\xi^{\mathrm{T}},v^{\mathrm{T}}]^{\mathrm{T}}$ as a state and $\lambda$ as a new input, this system falls into the standard CBF framework because the input dependence of $h$ can be eliminated. The CBF obtained by this augmentation is called the integral CBF (I-CBF) (Ames et al., 2021). From a different perspective, the adoption of I-CBF allows the model to have an input dependence on the function $\Xi$. This motivated us to solve a problem
$\begin{cases}\lambda^{\ast}=&\mathop{\mathrm{arg~{}min}}\limits_{\lambda}\frac{\mathrm{d}}{\mathrm{d}t}\|v-k(x)\|^{2}+\beta\|\lambda\|^{2}\\ \mathrm{s.t.}&\frac{\partial h}{\partial\xi}(f(\xi)+g(\xi)v)+\frac{\partial h}{\partial v}\lambda\leq\alpha(-h(\xi,v))\end{cases},$ (34)
with a positive constant $\beta$, to determine $\lambda$ as $\lambda=\lambda^{*}$. The first term aims at the quick convergence of $v$ to $k(\xi)$, and the second term is added to regularize the time-derivative of the input $v$. Although the input constraint $v\in V$ originally considered in (30) disappears in the above problem, it can be treated by combining the vector function
$h_{v}(\xi,v):=\begin{bmatrix}v-\Psi(\bar{u},\Phi^{-1}(\xi))\\ \Psi(\underline{u},\Phi^{-1}(\xi))-v\end{bmatrix}$ (35)
to $h(\xi,v)$ for handling the input constraint (1). In summary, this approach can handle the constraints (1) and (2) by setting $h(\xi,v)=[h_{y}(\xi,v)^{\mathrm{T}},h_{v}(\xi,v)^{\mathrm{T}}]^{\mathrm{T}}$.
The function to be minimized in (34) is quadratic with respect to $\lambda$, and it is given by
$F(\lambda)$ $:=\frac{\mathrm{d}}{\mathrm{d}t}\|v-k(\xi)\|^{2}+\beta\|\lambda\|^{2}$ (36)
$=\beta\|\lambda\|^{2}+2(v-k(\xi))^{\mathrm{T}}\lambda-2(v-k(\xi))^{\mathrm{T}}\frac{\partial k(\xi)}{\partial\xi}(f(\xi)+g(\xi)v).$
In addition, any constraint expressed in the form $h(\xi,v)\leq 0$ is reduced to the affine constraint for $\lambda$ in the I-CBF approach. Thus, the above optimization problem is a QP and can be solved quickly using many ready-made solvers. Further, the EL model allows functions $h(\xi,v)$ corresponding to (1) and (2) to be convex to $v$ for any fixed $\xi$. This guarantees that $\xi,v$ at the equilibrium point of this system is the solution to the problem $\mathop{\mathrm{minimize}}\limits_{v}\|v-k(\xi)\|^{2}\ \mathrm{s.t.}\ h(\xi,v)\leq 0$ (See Appendix 3).
Note that, in practice, the above controller is implemented in a discrete-time manner like $v_{k+1}=v_{k}+\lambda_{k}\Delta t$, where $\Delta t$ is a sampling period.
@@ 4 Experimental Evaluation
Table 1: Configuration for model learning
Item Setting
$N_{\Phi}$ (Number of layers for $\Phi$) 6
Number of neurons for $\Phi$ 3 (for all layers)
$N_{\Psi}$ (Number of layers for $\Psi$) 6
Number of neurons for $\Psi$ 3 (for all layers)
$N_{\Xi}$ (Number of layers for $\Xi$) 2
Number of neurons for $\Xi^{*1}$ 30 (for all layers)
Activation functions for $\Xi$ $\varphi_{\Xi}^{(1)}(x)={\rm softplus}(x)$, $\varphi_{\Xi}^{(2)}(x)=x$, $\varphi_{\Xi d}^{(1)}(x)={\rm softplus}(x)$
$N_{d}$ (Number of layers for FNN∗2) 2
Number of neurons for FNN∗2 30 (for all layers)
Activation functions for FNN∗2 $\varphi^{(1)}(x)=\max(0.1\cdot x,x)$, $\varphi^{(2)}(x)=x$
Weight matrix $Q_{e}$ diag([1,1,1,1,1])
$N_{s}$ (Number of samples) 831,488
*1: The setting is applied for $\varphi_{\Xi}^{(i)},\varphi_{\Xi d}^{(i)}\ (i=1,\ldots,N_{\Xi})$. *2: The setting is applied for $\varphi_{\Phi d}^{(i)},\varphi_{\Psi d}^{(i)},\varphi_{\Psi x}^{(i)},\varphi_{s}^{(i)}\ (i=1,\ldots,N_{d})$ in Fig. 2. Each layers implicitly includes weights and biases of appropriate size to be learnt.
The learning results are compared with those of the S–HW model (Moriyasu et al., 2022), EL model A (output $y$ is independent of the internal input $v$), and EL model B (output $y$ depends on the internal input $v$). The model learning settings were determined by trial and error and are listed in Table 1. The response results $\left\{x,y\right\}$ of each model to the input data $\left\{u,d\right\}$ are shown in Fig. 8, and the comparison of the root mean squared error (RMSE) is listed in Table 8. Fig. 8 shows that the response of the EL model (A, B) is qualitatively better fitted to validation data than that of the S–HW model. Table 8 shows that there is a large difference in the coefficient of determination of the state $x$ between the S–HW and EL models (A, B). The results confirm that the modeling accuracy can be improved by adding the dependency of the internal input $v$ on the output $y$. The target system has complex multi-physics, and therefore, whether it is EL or not has not been verified by physical modeling. However, since some outputs are significantly more accurate with EL than with S–HW, it can be inferred that the target system contains characteristics that cannot be inherently represented with S–HW.
Figure 6: Results of model prediction
Figure 7: Control result of input $u$
Table 2: Root mean squared error
S–HW EL(A) EL(B)
$x_{1}$ 17.72 16.06 15.49
$x_{2}$ 0.245 0.189 0.140
$x_{3}$ 1.218 0.546 0.573
$y_{1}$ 2.668 2.600 2.140
$y_{2}$ 0.505 0.404 0.377
Figure 8: Control result of state $x$ and output $y$
Next, we show the control simulation results. The learned EL model (B) is used as the control model and also as the system to be controlled; i.e., assuming no modeling error; since the main focus is on the nominal behavior of the proposed methods. We designed two types of controllers using the methods described in 3.1 and 3.2. That is, the former one is LQR (26) and the later solves QP (34) to calculate the control input online using LQR (26) as the reference function $k(\xi)$. Weight matrices are set to be $Q={\rm diag}([0.1,0.1,0.1]),R={\rm diag}([10,10,10])$ and regularization parameter is $\beta=10$. These were determined by trial and error. The sampling time is 0.001 s, the input and output constraints represented by (1) and (2) are $0\leq u_{i}\leq 100\,(i=1,\ldots,3)$ and $y_{1}\leq 80,y_{2}\leq 6.5$. In addition, the class $\mathcal{K}_{\infty}$ functions $\alpha_{i},\alpha_{j}\,(i=1,\ldots 3,j=1,2)$ in (34) are quadratic. The numerical results are shown in Figs. 8 and 8. In these figures, each result and color scheme are shown below.
  • Reference $x_{\mathrm{d}}$ for the state $x$ (black solid),
  • Bounds for the input $u$ and output $y$ (black dashed),
@@
We can see that QP realized the regulation control of state $x$ to the reference $x_{\mathrm{d}}$ while satisfying constraints (1) and (2). For input $u$, the control law based on (26) (blue line) fails to satisfy the constraint (1) (black dashed line), whereas the control law based on QP (green line) is within the constraints. For state $x$, the results of the QP deviate slightly from the reference $x_{\mathrm{d}}$ because there is no input $u$ such that $v=\Psi(u,x,\bar{d})$ satisfies constraints (1) and (2) under $x=x_{\mathrm{d}}$. In other words, from (24), it is possible to obtain $\xi_{\mathrm{d}}$ and $v_{\mathrm{d}}$ corresponding to $x_{\mathrm{d}}$; however, $u,y$ obtained by $\Psi$ in (10) and $\Xi$ in (13) do not satisfy the constraints (1), (2). Fig. 8 suggests that our proposed method automatically find a suitable steady input $u$ that achieves good regulation performance for the target $x_{\mathrm{d}}$ while satisfying constraints (1) and (2), which is the purpose of this study. The computation time of the proposed contoller based on QP was 0.396 ms for each control period on Intel Core i7-1265U processor, showing sufficiently fast computation. We used QP-KWIK solver (Schmid and Biegler, 1994) to solve QP (34).
In practice, we cannot assume that there is no modeling error and the target system have to be replaced into an actual plant or a high-fidelity simulator from which learning data have been obtained. In cases involving modeling errors, it is effective to configure a servo system that includes a disturbance observer and/or an integrator, but this is not the main focus of this paper and is omitted here.